# Comparison of Generative Learning Methods for Turbulence Modeling

### v1

@@ Abstract
Numerical simulations of turbulent flows present significant challenges in fluid dynamics due to their complexity and high computational cost. High resolution techniques such as Direct Numerical Simulation (DNS) and Large Eddy Simulation (LES) are generally not computationally affordable, particularly for technologically relevant problems. Recent advances in machine learning, specifically in generative probabilistic models, offer promising alternatives for turbulence modeling. This paper investigates the application of three generative models — Variational Autoencoders (VAE), Deep Convolutional Generative Adversarial Networks (DCGAN), and Denoising Diffusion Probabilistic Models (DDPM) — in simulating a 2D Kármán vortex street around a fixed cylinder. Training data was obtained by means of LES. We evaluate each model’s ability to capture the statistical properties and spatial structures of the turbulent flow. Our results demonstrate that DDPM and DCGAN effectively replicate the flow distribution, highlighting their potential as efficient and accurate tools for turbulence modeling. We find a strong argument for DCGAN, as although they are more difficult to train (due to problems such as mode collapse), they gave the fastest inference and training time, require less data to train compared to VAE and DDPM, and provide the results most closely aligned with the input stream. In contrast, VAE train quickly (and can generate samples quickly) but do not produce adequate results, and DDPM, whilst effective, is significantly slower at both inference and training time.
@@
Deep convolutional generative adversarial networks , Denoising diffusion probabilistic models , Variational autoencoders , Turbulence modeling , Kármán vortex street
@@ 1 Introduction
Turbulent flows are notoriously difficult to model. The structures involved can be found across a wide range of temporal and spatial scales, and the high degree of non-linearity as well as sensitivity to the initial conditions make this an especially challenging problem [1]. In particular, even in the presence of statistical or geometrical symmetries it is not possible to reduce the dimensionality of the problem (e.g. from three- to two- dimensional) in numerical simulations.
@@
Furthermore, these kind of flows play a central role in the field of fluid dynamics, and therefore in diverse fields such as aerospace [2], astrophysics [3], quantum mechanics [4], and even in immune biology [5], since almost all real world flows exhibit turbulence of some kind. Despite the best efforts and decades of research from physicists and mathematicians, a general analytic solution to the governing (Navier-Stokes) equations remains elusive, requiring a variety of computational methods to obtain even numerical solutions. Scale-resolving techniques, such as DNS and LES, strive to resolve the entire spectrum of length and time scales present in the flow, or at least the energetically most significant part (as is the case for LES), whereas conventional averaging approaches (such as Reynolds-averaged Navier-Stokes, or RANS) instead attempt to fit a statistical model. Whilst the former approaches provide the highest possible modelling accuracy, the latter are affected by inevitable loss of information, with non-generalisable closures, strongly affected by flow typology and boundary conditions. Even in techniques such as RANS, which considerably reduce the computational cost of complex simulations, they remain prohibitively high if a large number of computations must be carried out in relatively short time, such as in rapid prototyping and design optimisation loops [6]. This makes the use of machine learning, especially generative probabilistic AI extremely promising, particularly when only statistical distributions resulting from stochastic initial conditions are required [7, 8].
@@
The availability of large amounts of data, helped by a recent increase in computing power and specialised machine learning hardware, as well as the development of new, bleeding edge model architectures, allow the issue of large computation cost in machine learning to be addressed conveniently, although these are still not remotely comparable to the costs induced by full numerical simulations.
@@
The authors of [9] and [10] pioneered the use of deep neural networks (DNNs) to determine the model constants of nonlinear algebraic vortex viscosity models, significantly improving the prediction of anisotropic turbulence effects.
@@
While early work used machine learning mainly to improve the prediction quality of existing models, also known as ML-augmented turbulence modeling [11, 12, 13, 14, 15, 16, 17, 18], recent publications propose that turbulent flow can be modeled directly by generative models.
@@
In addition to modeling entire flow fields, the applications of generative models range from improving the resolution of flow fields [19], to reconstructing incomplete (gappy) flow fields using the in-painting technique [20], to uncertainty quantification to assess the variability in predictions due to uncertainties in initial or boundary conditions, or indeed in the flow model itself [21, 22].
@@
New generative models are one of the driving forces behind the growing application of generative learning to turbulence modeling. Kingma et. al. [23] introduced the Variational Autoencoder (VAE), a first generative model that provides stochastic variational inference and a learning algorithm that scales to large data sets. Shortly thereafter, Goodfellow et al. [24] proposed a generative adversarial network (GAN) that uses a minimax game to train an optimal generator, providing a paradigm shift in generative learning. Recently, the diffusion probabilistic models (DDPM) was introduced by Ho et. al. [25] as another powerful generative model, which is based on the principle of diffusion models [26] where the distribution of the data is learned by an iterative Markov chain process.
@@
The beauty of all these generative models is that the production of the data at inference time is very inexpensive once the models are trained. This allows for a large volume of data to be created, producing a representation of the target distribution that would be simply impractical to obtain with traditional methods.
@@
The question that remains is whether the results of the generative models are of reasonable quality and which generative model is best suited for modelling 2D turbulence, and in particular, a flow field around a cylinder. We compare three generative models: VAE, Deep Convolutional GAN (DCGAN) [27], and DDPM, and discuss their capabilities and limitations in terms of visual quality, physics-based metrics and computational cost.
@@
Our previous work [28] already showed that GAN-synthesized turbulence match LES-flow excellently. We show in this paper that the DDPM’s results are competitive with GAN, but require significantly more training data, supporting the findings of [29], and are computationally intensive in terms of both training and sample generation. The average inference speed per sample of our DDPM model is $36.3s$ compared to that of $0.001s$ by the GAN, a speedup factor of more than $1000$. The VAE can generate at $0.0003$ samples per second, but suffers from physical and visual inaccuracies.
@@
The remainder of the paper is structured as follows. Section 2 gives a review on key studies in the field using VAE, GAN or DDPM for applications of turbulence modeling. In section 3 we describe our specific implementation of the VAE, GAN, and DDPM along with a summary of the theoretical foundations of these models. Section 4 offers a brief description of the dataset used in this study. In section 5, we will discuss each model’s performance using physics-based evaluation metrics to assess their effectiveness and examine the associated computational costs. Finally, section 6 will summarize the main findings, discuss their broader implications, and suggest potential avenues for future research.
@@ 2 Related work
The review by Zhang et al. [30] gives an overview of the historical development of turbulence modeling approaches, starting with RANS and moving on to LES and modern deep learning methods. In the field of classical simulation methods, RANS can provide solutions at a comparably lower cost, but also paying the penalty of a reduced accuracy. Machine learning approaches represent a promising solution to overcome the problem of high computational cost without losing the details of turbulent structures [31, 32, 33]. In this section we give a broad overview of contributions using VAE, GAN and, DDPM.
@@
Over the last few years, modified and extended versions of the Variational Autoencoder have been developed to solve problems in a variety of scientific fields [34, 35, 36, 37, 38]. Advanced VAE frameworks have also been applied to turbulence generation. The authors of [39] presented a semi-conditional VAE (SCVAE) to reconstruct nonlinear flow from spatially sparse observations, which also allows, due to the probabilistic reconstruction, uncertainty quantification of the prediction. In [40], a two-stage approach is proposed where low-dimensional dynamics reconstructed by an autoencoder are enhanced by another high-resolution neural network. The work of [41] and [42] goes even further and uses a hybrid approach to model turbulence, combining an autoencoder with a multilayer perceptron network or a GAN to predict steady flow fields around supercritical airfoils and nonlinear fluid flows in varying parameterised space.
@@
Among the new approaches to turbulence modeling, many make use of GAN technology. In our previous work, we introduced GAN as a mathematically well-founded approach for synthetic modeling of turbulent flows [28]. Pioneering this field of research, King et al. [43, 44] showed that GANs are capable of generating syntheses of 2D flow fields after having been previously trained on direct numerical simulation (DNS) data. The reproductions even satisfied statistical constraints of turbulent flows, such as the Kolmogorov - 5/3 law and the small-scale intermittency of turbulence. Using an unsupervised trained combination of a GAN and a recurrent neural network (RNN), Kim and Lee synthesized boundary conditions for turbulent flow [45] or generated stationary DNS flow fields [46].
With the use of more advanced GAN frameworks such as conditional GAN (cGAN), turbulence can be predicted and controlled [47], even only at specific local points [48]. Li et. al. investigated the task of inferring a velocity component from the measurement of a second one for a rotating turbulent flow. To do this, a GAN with context encoders takes the simulated data as input and adds a second loss term that measures the point-to-point distance between the GAN output and the ground truth configuration.
Another application of GAN is the large field of super-resolution reconstruction of turbulent flows. These methods can be used to synthetically scale up low-resolution or noisy flow fields [19, 49, 50, 51, 52, 53, 54, 55]. For these problems, the so-called super-resolution GAN (SRGAN) [56] or enhanced SRGAN (ESRGAN) [57] are used, which can be conditioned by additional physical information [58]. Also the high-fidelity reconstruction of 2D damaged turbulent fields was addressed with GAN [20, 59].
Finally, the generalization capabilities of GAN are of interest and have been investigated. In our previous work, we demonstrated the generalization capabilities of conditional deep convolutional turbulence generators when geometric changes occur in the flow configuration [60]. By changing parameters in the numerical setup, such as the Reynolds number, Nista et. al. investigated the generalization capabilities of their proposed SRGAN [61]. The authors of [62] combined high and low resolution flows to improve the generalization capability for a physics-informed super-resolution GAN (SRGAN).
@@
The most recent of the generative models considered in our work are the Denoising Diffusion Probabilistic Models [25] based on the idea of Sohl et. al. [26]. In the field of turbulence modeling, diffusion models are gaining popularity, especially for probabilistic spatiotemporal forecasting. For example, Gao et. al. [63] implemented a Bayesian conditional diffusion model for versatile spatiotemporal turbulence generation. Whereas autoregressive learning with DDPM is often used to solve the task of probabilistic spatiotemporal prediction, cf. [64, 65, 66], Rühling et. al. [67] introduced a dynamically informed diffusion model, which adapts the model to the dynamic nature of the data in order to achieve long-range prediction at inference time. Another application of DDPM in turbulence modeling is the reconstruction of high-resolution turbulent flow fields from low-resolution flow data. For example, Qi et. al. [68] combined a conditional DDPM with an enhanced residual network, and Sardar et. al. [69] developed a preprocessing to decompose flow fields into high and low wavenumber components to learn a conditional DDPM on this data. Furthermore, DDPM were also investigated for training uncertainty-aware surrogate models for simulating turbulence as flows around differently shaped airfoils [70].
@@ 3.1 Variational Autoencoders (VAE)
Figure 1: Comparison of the classical [71] (top) and the variational [72] autoencoder (bottom) architectures. In the case of classical AE, the encoder $f_{\theta}$ receives as input a real-world image $x$ whose most important features are encoded in the latent space $\mathcal{Z}^{\prime}$. Sending a point of this latent space back to the decoder $g_{\phi}$ should result in a reconstructed image $x^{\prime}\approx x$. In contrast, in the case of the VAE, the real images have a chosen distribution, making the encoder and decoder probabilistic. The encoder learns the distribution parameters of the input images, and latent samples are drawn during training by the re-parameterization trick, making it possible to compute the gradients for $\mu$ and $\sigma$. Due to the stochasticity of the probabilistic decoder’s input, the resulting images are generated rather than reconstructed, as they may differ from the original images.
@@
An autoencoder consists of two components - an encoder $f:\mathcal{X}\rightarrow\mathcal{Z}^{\prime}$ and a decoder $g:\mathcal{Z}^{\prime}\rightarrow\mathcal{X}^{\prime}$ with $\mathcal{X}$ the space of real-world images, $\mathcal{Z}^{\prime}$ the latent space and $\mathcal{X}^{\prime}$ the space of reconstructed images. In machine learning, both functions are represented by neural networks, and in our work specifically we use deep convolutional neural networks (DCNN), which are parameterized by weights $\phi$ and $\theta$. Due to the way an AE is implemented and the definition of its loss function, it can be trained end-to-end. The goal of training an AE is to learn a decoding function that is able to reconstruct an original image from a point in the encoded latent space such that $x^{\prime}=g_{\theta}(z^{\prime})=g_{\theta}(f_{\phi}(x))\approx x$ with $z^{\prime}\in\mathcal{Z}^{\prime},x\in\mathcal{X}$ and $x^{\prime}\in\mathcal{X}^{\prime}$ (see figure 1). Once trained, the process of generating samples, which we also refer to as inference, is a simple matter of picking a point in the latent space and passing it through the decoding function. Typically, the latent space in which important features are embedded is chosen to be small. This property makes AE also a dimensionality reduction method and can be seen as a generalization of principal component analysis (PCA) [73, 71].
@@
In recent years, more advanced variants of the AE have been developed [74], including the variational autoencoder (VAE) [72]. In contrast to a classical AE, the real-world input images $x\in\mathcal{X}$ are drawn from a chosen probability distribution $p_{\theta}(x)$ parameterized by the network weights $\theta$, making the encoder and decoder probabilistic and the latent space a latent distribution encoding the distribution parameters. In general, this underlying distribution is chosen to be the Gaussian $\mathcal{N}(x;\mu,\sigma)$. Through the application of the powerful mathematical framework of Bayesian probability [75, 76], the VAE can be interpreted as a model of the joint distribution $p_{\theta}(x,z)=p_{\theta}(x|z)p_{\theta}(z)$ of the real-world images $x\in\mathcal{X}$ and the samples of the latent distribution $z\in\zeta$, where the goal is to compute the posterior

### v2

# Comparison of Generative Learning Methods as Turbulence Surrogates

@@ Abstract
Numerical simulations of turbulent flows present significant challenges in fluid dynamics due to their complexity and high computational cost. High resolution techniques such as Direct Numerical Simulation (DNS) and Large Eddy Simulation (LES) are generally not computationally affordable, particularly for technologically relevant problems. Recent advances in machine learning, specifically in generative probabilistic models, offer promising alternatives as surrogates for turbulence. This paper investigates the application of three generative models — Variational Autoencoders (VAE), Deep Convolutional Generative Adversarial Networks (DCGAN), and Denoising Diffusion Probabilistic Models (DDPM) - in simulating a von Kármán vortex street around a fixed cylinder projected into 2D, as well as a real-world experimental dataset of the wake flow of a cylinder array. Training data was obtained by means of LES in the simulated case and Particle Image Velocimetry (PIV) in the experimental case. We evaluate each model’s ability to capture the statistical properties and spatial structures of the turbulent flow. Our results demonstrate that DDPM and DCGAN effectively replicate all flow distributions, highlighting their potential as efficient and accurate tools for turbulence surrogacy. We find a strong argument for DCGAN, as although they are more difficult to train (due to problems such as mode collapse), they show the fastest inference and training time, require less data to train compared to VAE and DDPM, and provide the results most closely aligned with the input stream. In contrast, VAE train quickly (and can generate samples quickly) but do not produce adequate results, and DDPM, whilst effective, are significantly slower at both, inference and training time.
@@
Deep convolutional generative adversarial networks , Denoising diffusion probabilistic models , Variational autoencoders , Turbulence surrogacy , von Kármán vortex street
@@ 1 Introduction
Turbulent flows are notoriously difficult to model. The structures involved can be found across a wide range of temporal and spatial scales, and the high degree of non-linearity as well as sensitivity to the initial conditions make this an especially challenging problem [frisch1995turbulence].
@@
Furthermore, these kind of flows play a central role in the field of fluid dynamics, and therefore in diverse fields such as aerospace [Spalart_Venkatakrishnan_2016], astrophysics [stars_turbulence], quantum mechanics [Barenghi_2014], and even in immune system biology [cancer_turbulence], since almost all real world flows exhibit turbulence of some kind. Despite the best efforts and decades of research from physicists and mathematicians, a general analytic solution to the governing (Navier-Stokes) equations remains elusive, requiring a variety of computational methods to obtain even numerical solutions. Scale-resolving techniques, such as DNS and LES, strive to resolve the entire spectrum of length and time scales present in the flow, or at least the energetically most significant part (as is the case for LES), whereas conventional averaging approaches (such as Reynolds-averaged Navier-Stokes, or RANS) instead attempt to fit a statistical model. Whilst the former approaches provide the highest possible modeling accuracy, the latter are affected by inevitable loss of information, with non-generalizable closures, strongly affected by flow typology and boundary conditions. Even in techniques such as RANS, which considerably reduce the computational cost of complex simulations, they remain prohibitively high if a large number of computations must be carried out in relatively short time, such as in rapid prototyping and design optimization loops [gopalan2013unified]. This makes the use of machine learning, especially generative probabilistic AI extremely promising, particularly when only statistical distributions resulting from stochastic initial conditions are required [lumley2007stochastic, olson2001turbulence].
@@
The availability of large amounts of data, helped by a recent increase in computing power and specialized machine learning hardware, as well as the development of new, bleeding edge model architectures, allow the issue of large computation cost in machine learning to be addressed conveniently, although these are still not remotely comparable to the costs induced by full numerical simulations.
@@
Ling et al. [Ling:2016] and Jiang et al. [Jiang:2020] pioneered the use of deep neural networks (DNNs) to determine the model constants of nonlinear algebraic vortex viscosity models, significantly improving the prediction of anisotropic turbulence effects.
@@
While early work used machine learning mainly to improve the prediction quality of existing models, also known as ML-augmented turbulence modeling [Cheung:2011, Edeling:2014a, Edeling:2014b, Weatheritt:2016, Weatheritt:2017, Zhang:2018a, Zhang:2018b, Zhao:2020], recent publications propose that turbulent flow can be modeled directly by generative models.
@@
In addition to modeling entire flow fields, the applications of generative models range from improving the resolution of flow fields [deng2019super], to reconstructing incomplete (gappy) flow fields using the in-painting technique [zheng2024high], to uncertainty quantification to assess the variability in predictions due to uncertainties in initial or boundary conditions, or indeed in the flow model itself [cheng2023uncertainty, goswami2023uncertainty]. Also Echo State Networks were recently used to spatially and/or temporally predict turbulent flow fields based on temporally resolved and spatially under-resolved experimental data [Sharifi_Ghazijahani_2025_eaai, Sharifi_Ghazijahani_2024_scirep, Ghazijahani_2024_MLST].
@@
New generative models are one of the driving forces behind the growing application of generative learning to turbulence modeling. Kingma et al. [kingma2013] introduced the Variational Autoencoder (VAE), a first generative model that provides stochastic variational inference and a learning algorithm that scales to large data sets. Shortly thereafter, Goodfellow et al. [vanilla_gan] proposed a generative adversarial network (GAN) that uses a minmax game to train an optimal generator, providing a paradigm shift in generative learning. Recently, the diffusion probabilistic models (DDPM) was introduced by Ho et al. [ho2020denoising] as another powerful generative model, which is based on the principle of diffusion models [sohldickstein2015deep] where the distribution of the data is learned by an iterative Markov chain process. Recently, Teutsch et al. [Teutsch_2025] showed that convolutional autoencoder-based reduced-order models can not only be used for modeling turbulence but provide interpretable features that further help to understand the flow physics.
@@
The advantage of these generative models is that the production of the data at inference time is very inexpensive once the models are trained. This allows for a large volume of data to be created, producing a representation of the target distribution that would be simply impractical to obtain with traditional methods.
@@
The question that remains is whether the results of the generative models are of reasonable quality and which generative model is best suited for modeling turbulence, and in particular, a flow field around a cylinder and in the wake of a cylinder array. We compare three generative models: VAE, Deep Convolutional GAN (DCGAN) [dcgan], and DDPM, and discuss their capabilities and limitations in terms of visual quality, physics-based metrics and computational cost.
@@
Our previous work [drygala2022generative] already showed that GAN-synthesized turbulence match LES-flow excellently. We show in this paper that the DDPM’s results are competitive with GAN, but require significantly more training data, supporting the findings of [wang2024patch], and are computationally intensive in terms of both training and sample generation. The average inference speed per sample of our DDPM model is $36.3s$ compared to that of $0.001s$ by the GAN, a speedup factor of more than $1,000$. The VAE can generate at $0.0003$ samples per second, but suffers from physical and visual inaccuracies.
With this work we place at least one of the hyperparameters of the machine learning approach for turbulence (namely, the model architecture) on scientific footing. We also note that we are one of the first to train generative machine learning model on a high-fidelity experimental dataset. In particular, we only check to see if the models can accurately reproduce the statistics of the projected flow. While the majority of the ML community are moving towards diffusion and other flow-matching based models due to the success of models like StableDiffusion [latentdiffusionmodel2022] in other domains, this article finds in favor of GANs when applied to turbulence.
@@
The remainder of the paper is structured as follows. Section˜2 gives a review on key studies in the field using VAE, GAN or DDPM for applications of turbulence modeling. In section˜3 we describe our specific implementation of the VAE, GAN, and DDPM along with a summary of the theoretical foundations of these models. Section˜4 offers a brief description of the experimental and numerical datasets used in this study. In section˜5, we will discuss each model’s performance using physics-based evaluation metrics to assess their effectiveness and examine the associated computational costs. Finally, section˜6 will summarize the main findings, discuss their broader implications, and suggest potential avenues for future research.
@@ 2 Related work
Turbulence modeling has recently seen the rise of data-driven and machine-learning-based methods that complement classical approaches such as RANS and LES, with reviews of both the historical development and contemporary ML techniques provided in the recent literature [duraisamy2019turbulence, zhangreview2023, fang2025breakthroughs]. In the field of classical simulation methods, RANS can provide solutions at a comparably lower cost, but also paying the penalty of a reduced accuracy. Machine learning approaches represent a promising solution to overcome the problem of high computational cost without losing the details of turbulent structures [wang_2017, Maulik_2018, shankar2023differentiableturbulenceii]. In this section we give a broad overview of contributions using VAE, GAN and, DDPM.
@@
Over the last few years, modified and extended versions of the Variational Autoencoder have been developed to solve problems in a variety of scientific fields [higgins2017betavae, DBLP:journals/corr/abs-1903-05789, ghosh2020variationaldeterministicautoencoders, NIPS2016_ddeebdee, oord2018neuraldiscreterepresentationlearning]. Advanced VAE frameworks have also been applied to turbulence generation. Gunderson et al. [gundersen2021scvae] presented a semi-conditional VAE (SCVAE) to reconstruct nonlinear flow from spatially sparse observations, which also allows, due to the probabilistic reconstruction, uncertainty quantification of the prediction. In 2020, Agostini [agostini2020aelow], a two-stage approach is proposed where low-dimensional dynamics reconstructed by an autoencoder are enhanced by another high-resolution neural network. The work of Wang et al. [wang2021flow] and Cheng et al. [CHENG2020113375] goes even further and uses a hybrid approach to model turbulence, combining an autoencoder with a multilayer perceptron network or a GAN to predict steady flow fields around supercritical airfoils and nonlinear fluid flows in varying parameterized space. As often the availability of data is problematic, Teutsch et al. [Teutsch2025_aplml] presented a pipeline that determines useful data augmentation strategies for different turbulent flows.
@@
Among the new approaches to turbulence modeling, many make use of GAN technology. In our previous work, we introduced GAN as a mathematically well-founded approach for synthetic modeling of turbulent flows [drygala2022generative]. Pioneering this field of research, King et al. [King:2017, King:2018] showed that GANs are capable of generating syntheses of 2D flow fields after having been previously trained on direct numerical simulation (DNS) data. The reproductions even satisfied statistical constraints of turbulent flows, such as the Kolmogorov - 5/3 law and the small-scale intermittency of turbulence. Using an unsupervised trained combination of a GAN and a recurrent neural network (RNN), Kim and Lee synthesized boundary conditions for turbulent flow [Kim:2020] or generated stationary DNS flow fields [kim2021unsupervised].
With the use of more advanced GAN frameworks such as conditional GAN (cGAN), turbulence can be predicted and controlled [Kim_Kim_Lee_2024], even only at specific local points [yan2023local]. [atmos15010060] investigated the task of inferring a velocity component from the measurement of a second one for a rotating turbulent flow. To do this, a GAN with context encoders takes the simulated data as input and adds a second loss term that measures the point-to-point distance between the GAN output and the ground truth configuration.
Another application of GAN is the large field of super-resolution reconstruction of turbulent flows. These methods can be used to synthetically scale up low-resolution or noisy flow fields [deng2019super, Fukami:2019, Fukami:2020, Liu:2020, Xie:2017, Werhahn:2019, youssif2021high, youssif2022superres]. For these problems, the so-called super-resolution GAN (SRGAN) [ledig2017photo] or enhanced SRGAN (ESRGAN) [wang2018esrgan] are used, which can be conditioned by additional physical information [subramaniam2020turbulence]. Also the high-fidelity reconstruction of 2D damaged turbulent fields was addressed with GAN [zheng2024high, Li_Buzzicotti_Biferale_Bonaccorso_Chen_Wan_2023].
Finally, the generalization capabilities of GAN are of interest and have been investigated. In our previous work, we demonstrated the generalization capabilities of conditional deep convolutional turbulence generators when geometric changes occur in the flow configuration [drygala2023generalization]. By changing parameters in the numerical setup, such as the Reynolds number, Nista et al. investigated the generalization capabilities of their proposed SRGAN [nista2023investigation]. Bode et al. [bode2021using] combined high and low resolution flows to improve the generalization capability for a physics-informed super-resolution GAN (SRGAN).
@@
The most recent of the generative models considered in our work are the Denoising Diffusion Probabilistic Models [ho2020denoising] based on the idea of Sohl et al. [sohldickstein2015deep]. In the field of turbulence modeling, diffusion models are gaining popularity, especially for probabilistic spatiotemporal forecasting. For example, Gao et al. [gao2024bayesian] implemented a Bayesian conditional diffusion model for versatile spatiotemporal turbulence generation. Whereas autoregressive learning with DDPM is often used to solve the task of probabilistic spatiotemporal prediction, cf. [shu2024zero, kohl2024benchmarking, yang2023diffusion], Rühling et al. [NEURIPS2023_8df90a14] introduced a dynamically informed diffusion model, which adapts the model to the dynamic nature of the data in order to achieve long-range prediction at inference time. Another application of DDPM in turbulence modeling is the reconstruction of high-resolution turbulent flow fields from low-resolution flow data. For example, Qi et al. [Qi2024Combined] combined a conditional DDPM with an enhanced residual network, and Sardar et al. [sardar2023spectrally] developed a preprocessing to decompose flow fields into high and low wavenumber components to learn a conditional DDPM on this data. Furthermore, DDPM were also investigated for training uncertainty-aware surrogate models for simulating turbulence as flows around differently shaped airfoils [liu2024uncertainty].
@@ 3.1 Variational Autoencoders (VAE)
Figure 1: Comparison of the classical [kramer1991nonlinear] (top) and the variational [kingma2013auto] autoencoder (bottom) architectures. In the case of classical AE, the encoder $f_{\theta}$ receives as input a real-world image $x$ whose most important features are encoded in the latent space $\mathcal{Z}^{\prime}$. Sending a point of this latent space back to the decoder $g_{\phi}$ should result in a reconstructed image $x^{\prime}\approx x$. In contrast, in the case of the VAE, the real-world images have a chosen distribution, making the encoder and decoder probabilistic. The encoder learns the distribution parameters of the input images, and latent samples are drawn during training by the re-parameterization trick, making it possible to compute the gradients for $\mu$ and $\sigma$. Due to the stochasticity of the probabilistic decoder’s input, the resulting images are generated rather than reconstructed, as they may differ from the original images.
@@
An autoencoder consists of two components - an encoder $f:\mathcal{X}\rightarrow\mathcal{Z}^{\prime}$ and a decoder $g:\mathcal{Z}^{\prime}\rightarrow\mathcal{X}^{\prime}$ with $\mathcal{X}$ the space of real-world images, $\mathcal{Z}^{\prime}$ the latent space and $\mathcal{X}^{\prime}$ the space of reconstructed images. In machine learning, both functions are represented by neural networks, and in our work specifically we use deep convolutional neural networks (DCNN), which are parameterized by weights $\phi$ and $\theta$. Due to the way an AE is implemented and the definition of its loss function, it can be trained end-to-end, meaning the encoder and decoder are optimized together in a single step, automatically learning useful internal representations without separate training. The goal of training an AE is to learn a decoding function that is able to reconstruct an original image from a point in the encoded latent space such that $x^{\prime}=g_{\theta}(z^{\prime})=g_{\theta}(f_{\phi}(x))\approx x$ with $z^{\prime}\in\mathcal{Z}^{\prime},x\in\mathcal{X}$ and $x^{\prime}\in\mathcal{X}^{\prime}$ (see figure˜1). Once trained, the process of generating samples, which we also refer to as inference, is a simple matter of picking a point in the latent space and passing it through the decoding function. Typically, the latent space in which important features are embedded is chosen to be small. This property makes AE also a dimensionality reduction method and can be seen as a generalization of principal component analysis (PCA) [ballard1987modular, kramer1991nonlinear].
@@
In recent years, more advanced variants of the AE have been developed [pinaya2020autoencoders], including the variational autoencoder (VAE) [kingma2013auto]. In contrast to a classical AE, the real-world input images $x\in\mathcal{X}$ are drawn from a chosen probability distribution $p_{\theta}(x)$ parameterized by the network weights $\theta$, making the encoder and decoder probabilistic and the latent space a latent distribution encoding the distribution parameters. In general, this underlying distribution is chosen to be the Gaussian $\mathcal{N}(x;\mu,\sigma)$. Through the application of the powerful mathematical framework of Bayesian probability [bayes1958essay, box2011bayesian], the VAE can be interpreted as a model of the joint distribution $p_{\theta}(x,z)=p_{\theta}(x|z)p_{\theta}(z)$ of the real-world images $x\in\mathcal{X}$ and the samples of the latent distribution $z\in\zeta$, where the goal is to compute the posterior